OpenStudy (anonymous):

find the derivative f(x)=-2x^4-5/3^x3-0.5x^2-4x-10

OpenStudy (lgbasallote):

\[\LARGE \frac{f'(x)}{g'(x)} = \frac{g(x) f'(x) - f(x) g'(x)}{g^2 (x)}\]

OpenStudy (anonymous):

but I thought you only do that when you have a f(x) and a g(x) like for example f(x)=3x^2+x and g(x)= 2x+1 Right ?

OpenStudy (lgbasallote):

your problem is \[\frac{-2x^4 - 5}{3x^3 - 0.5x^2 - 4x - 10}\] correct?

OpenStudy (lgbasallote):

if so i stand corrected. simplify it first

OpenStudy (anonymous):

In my problem where u have the 5 on top it was a fraction 5 over 3 x to the 3 power.

OpenStudy (fellowroot):